OpenStudy (anonymous):

Solve the system of equations using algebra. 5c -4t=8 14 +4t= 3c

OpenStudy (solomonzelman):

Addition or Substitution?

OpenStudy (anonymous):

elimination or substitution is what im guessing you mean.

OpenStudy (anonymous):

i use substitution.

OpenStudy (solomonzelman):

yes, my bad.

OpenStudy (solomonzelman):

Ok, then you have to set one of the equations as x=... or y... first.

OpenStudy (solomonzelman):

I would use elimination though, see how -4t and 4t are there already for you to add?

OpenStudy (anonymous):

the answers i got at the end were t= 15/11 and c= 40/11

OpenStudy (anonymous):

but i dont think they are right

OpenStudy (texaschic101):

if you use elimination, you need to set up the 2nd equation just like the first one by subtracting 3c and subtracting 14

OpenStudy (anonymous):

how would you do it withe elimination?

OpenStudy (solomonzelman):